OpenStudy (anonymous):

solve the equation 3(2^(x-3))=6^(x-1)

jhonyy9 (jhonyy9):

idea ? opinion ?

OpenStudy (anonymous):

idea

jhonyy9 (jhonyy9):

courage how we can solving the same equations when x is on the exponent ?

OpenStudy (kirbykirby):

^this mean's it's a candidate to use logarithms, since logarithms bring the exponent down

OpenStudy (anonymous):

how to eliminate the 3 in front of 2^(x-3) ?

OpenStudy (kirbykirby):

I'm trying to do it on paper. I'll get back to u in a few

OpenStudy (kirbykirby):

Ok got it. Give me some time to write it up

OpenStudy (anonymous):

ok

OpenStudy (kirbykirby):

\[3(2^{x-3})=6^{x-1}\] \(ln(3(2^{x-3}))=ln(6^{x-1})\) by taking ln on both sides \(ln(3)+ln(2^{x-3})=ln(6^{x-1})\) using log rule \(ln(a*b)=ln(a)+ln(b)\) \(ln(3)+(x-3)ln(2)=(x-1)ln(6)\) by using log rule \(ln(a^b)=b*ln(a)\) \(ln(3)+x*ln(2)-3ln(2)=x*ln(6)-ln(6)\) distributivity \(x*ln(6)-x*ln(2)=ln(3)-3ln(2)+ln(6)\) bring all x terms to one side \(x[ln(6)-ln(2)]=ln(3)-3ln(2)+ln(6)\) factor out x \[x=\frac{ln(3)-3ln(2)+ln(6)}{ln(6)-ln(2)}\]

OpenStudy (kirbykirby):

\[3(2^{x-3})=6^{x-1}\] \(ln(3(2^{x-3}))=ln(6^{x-1})\) by taking ln on both sides \(ln(3)+ln(2^{x-3})=ln(6^{x-1})\) using log rule \(ln(a*b)=ln(a)+ln(b)\) \(ln(3)+(x-3)ln(2)=(x-1)ln(6)\) by using log rule \(ln(a^b)=b*ln(a)\) \(ln(3)+x*ln(2)-3ln(2)=x*ln(6)-ln(6)\) distributivity \(x*ln(6)-x*ln(2)=ln(3)-3ln(2)+ln(6)\) bring all x terms to one side \(x[ln(6)-ln(2)]=ln(3)-3ln(2)+ln(6)\) factor out x \[x=\frac{ln(3)-3ln(2)+ln(6)}{ln(6)-ln(2)}\]

OpenStudy (kirbykirby):

ok hopefully that's a bit more clear, my line got shifted down

OpenStudy (anonymous):

thanks

OpenStudy (kirbykirby):

in decimal it's about 0.73814